OpenStudy (anonymous):

write the equation of a line with the given properties, in slope intercept form slope= -5 through (-4,2)

OpenStudy (anonymous):

y = -5x - 18

OpenStudy (anonymous):

can you show me how u got that? because i got y=-5x+22

OpenStudy (anonymous):

Since you're given your slope and a point you can use the equation \[y-y_1=m(x-x_1)\], where m is your slope and \[x_1\] and \[y_1\] are your coordinates of your point. So subbing what you have into this equation you get \[y-2=-5(x-(-4))\] Bringing the -2 across gives you: \[y-2=-5x-20\] Bring the 2 across: \[y=-5x-18\]

OpenStudy (anonymous):

y-y1=m(x-x) y-2=-5(x-(-4) y+2=-5x+20 y-2=-5x+02 y =-5x+22

OpenStudy (anonymous):

oo okay- it's a negative sign-- i see

OpenStudy (anonymous):

thanks